Suzan grabs five marbles at random from a bag containing four red marbles, three green ones, two white ones and one purple one. What are the probabilities of the following events expressing each as a fraction in the lowest term.

a. She has at least 2 green ones.

b. She does not have all the red ones.  
Answer #1

Ans:

There are 4 red marbles,3 green marbles,2 white marbles and 1 purple marble

We have to choose 5.

There are total possible 10C5 ways.

a)P(atleast 2 green)=1-P(0 green)-P(1 green)

=1-(7C5*3C0/10C5)-(3C1*7C4/10C5)

=1-(21/252)-(105/252)

=1-(126/252)

=1-(1/2)

=1/2

b)P(does not have all red)=1-P(all red)=1-(4C4*6C1)/10C5

=1-(6/252)

=246/252

=123/126